Find the general solution to y"+4y'+3y=x


1
Expert's answer
2021-10-03T17:34:45-0400

Characteristic equation: "r^2+4r+3=0\to r_1=-3,r_2=-1."

General solution of the homogeneous equation "y''+4y'+3y=0" is "y=C_1e^{-4x}+C_2e^{-x}" .

The particular solution might have the form: "y_p=Ax+b".

After substitution "y_p" into the equation we have: "4A+3Ax+3B=x".

So, "A=\frac{1}{3},B=-\frac{4}{9}."

General solution of our ODE: "y=C_1e^{-4x}+C_2e^{-x}+\frac{x}{3}-\frac{4}{9}."